OpenStudy (anonymous):

Two boats, A and B, are approaching a port. The distance between the boats is 12 miles. The bearings of the port with respect to boat A and boat B are N 60° W and N 35° E respectively. How far is boat A from the port?

OpenStudy (anonymous):

11.02 miles 9.87 miles 8.23 miles 4.54 miles

OpenStudy (dumbcow):

 use law of sines \[\frac{x}{\sin 55} = \frac{12}{\sin 95}\]

OpenStudy (anonymous):

So I solve that equation for x?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

Okay let me try it

OpenStudy (anonymous):

So B? @dumbcow